OpenStudy (anonymous):

Solve. x2 + 4x + 8 = 0 (2 points) x = –2 ± 2i x = –2 ± 4i x = –4 ± 2i x = –4 ± 4i

OpenStudy (whpalmer4):

Use the quadratic formula. a = 1, b = 4, c = 8. \[x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\] and \[\sqrt{-x} = i\sqrt{x}\]

OpenStudy (anonymous):

ive tried i don't understand it

OpenStudy (whpalmer4):

You just to plug in the numbers. \[x = \frac{-(4) \pm \sqrt{4^2-4(1)(8)}}{2(1)} = \frac{-4 \pm \sqrt{-16}}{2}\]

OpenStudy (whpalmer4):

What is \[\sqrt{-16}\]

OpenStudy (whpalmer4):

\[\sqrt{-16}=\sqrt{(-1)(16)}=\sqrt{-1}*\sqrt{16}\] \[\sqrt{-1} = i\]so\[\sqrt{-16}=i\sqrt{16}=4i\] Combine that with the expression for x we previously got.

OpenStudy (whpalmer4):

\[\frac{-4 \pm \sqrt{-16}}{2} = \frac{-4 \pm 4i}{2} \] and you should be able to figure out which of the answers is equal to that.